How do I do this problem?
The product of 2 integers is 162. Find the 2 integers.

Answer by checkley79(3341)

X*Y=162
ASSIGN A NUMBER FOR X BETWEEN 1 & 162 & SOLVE FOR Y.
EXAMPLE:
X=9
9Y=162
Y=162/9
Y=18 ANS.
PROOF:
9*18=162
162=162